# Language Models Fail to Introspect About Their Knowledge of Language

### v2

@@ Abstract
There has been recent interest in whether large language models (LLMs) can introspect about their own internal states. Such abilities would make LLMs more interpretable, and also validate the use of standard introspective methods in linguistics to evaluate grammatical knowledge in models (e.g., asking “Is this sentence grammatical?”). We systematically investigate emergent introspection across 21 open-source LLMs, in two domains where introspection is of theoretical interest: grammatical knowledge and word prediction. Crucially, in both domains, a model’s internal linguistic knowledge can be theoretically grounded in direct measurements of string probability. We then evaluate whether models’ responses to metalinguistic prompts faithfully reflect their internal knowledge. We propose a new measure of introspection: the degree to which a model’s prompted responses predict its own string probabilities, beyond what would be predicted by another model with nearly identical internal knowledge. While both metalinguistic prompting and probability comparisons lead to high task accuracy, we do not find evidence that LLMs have privileged “self-access”. Our findings complicate recent results suggesting that models can introspect, and add new evidence to the argument that prompted responses should not be conflated with models’ linguistic generalizations.
@@ 1 Introduction
Accordingly, there has been recent interest in whether large language models (LLMs) have abilities consistent with introspection, or metacognition more broadly (e.g., Thrush et al., 2024; Koo et al., 2024; Panickssery et al., 2024; Binder et al., 2025; Betley et al., 2025). Such abilities might be desirable for several reasons. From a safety perspective, an LLM that could report information about its internal states would be more interpretable and reliable (e.g., Binder et al., 2025; Betley et al., 2025). And from a scientific perspective, introspection would allow us to evaluate LLMs’ cognitive abilities using standard paradigms for measuring human linguistic knowledge, such as acceptability judgments (e.g., Dentella et al., 2023).
@@
Some recent studies have reported evidence that LLMs can introspect or demonstrate self-awareness of their own behaviors. For example, Binder et al. (2025) fine-tune LLMs to predict properties of their own behaviors in response to hypothetical prompts (e.g., “Suppose you were asked the following: {…} What is the second character of your output?”). The authors find that models are better at predicting their own hypothetical responses than other models fine-tuned on the same set of facts, concluding that models have privileged access to their own behaviors. Similarly, Betley et al. (2025) fine-tune LLMs to exhibit specific behaviors, such as outputting insecure code, and find that the resulting models can output explicit descriptions of these behaviors (e.g., “The code I write is insecure”).
@@
However, previous approaches to studying introspection face several limitations. First, Binder et al. found that models needed to be fine-tuned to show strong signs of introspection, raising the question of whether introspection can emerge in LLMs not specifically trained to predict their own outputs. Second, while Binder et al. find that models predict their own outputs better than other models do, there are subtleties involved in how to interpret these results in the light of major differences across models. If two models were trained on different datasets or have different inductive biases, and one model predicts its own outputs better than the other model’s outputs, this could be due to inherent differences between the two models, rather than genuine introspection. Third, studies of introspection in LLMs have used tasks that are quite different from the standard tasks used for metacognitive evaluation in humans, such as linguistic acceptability judgments. While this is not inherently problematic, it leaves open the question of whether metacognitive evaluation methods—which are prevalent in linguistics—are a valid way of measuring linguistic knowledge in LLMs (cf. Hu & Levy, 2023; Dentella et al., 2023; Hu et al., 2024).
@@
To foreshadow our results, we do not find compelling evidence of introspection in the LLMs tested. Instead, we find that models that are inherently more similar to each other—for example, differing only in their random seed initialization—have a stronger correlation between their metalinguistically- and directly-measured behaviors. In other words, for a pair of models and (including ), the correlation between ’s metalinguistic and ’s direct behaviors is predicted solely by the similarity between and , with no evidence for “privileged” self-access when . We conclude that prompted metalinguistic knowledge is real, but distinct from the knowledge of language used by models to assign probabilities to strings. Our findings complicate recent results suggesting that models can introspect, and also have important implications for researchers studying linguistic capabilities in models.
@@ 2 Overview of our approach
Our general approach is shown in Fig. 1. For each test item, we evaluate each model using two methods. The Direct method compares the log probabilities assigned to strings, in a way that theoretically corresponds to the model’s ground-truth knowledge state. For example, to evaluate a model’s word prediction abilities, we compare the probabilities assigned to two candidate words conditioned on a given prefix (LABEL:fig:introduction-inputs-prediction, top). The Meta method compares the log probabilities assigned to responses to metalinguistic prompts. For example, to evaluate word prediction, we might compare the probabilities assigned to the tokens “1” and “2” after a prompt like “Which word is a better continuation after [PREFIX], 1 or 2? 1: [WORD1], 2: [WORD2]. Respond with 1 or 2 as your answer.” (LABEL:fig:introduction-inputs-prediction, bottom). These two evaluation methods give us two quantities, each of which represents the model’s preference for one answer over another: and .
@@ 3.1 Methods
Under the Meta method, we measure the preference using 8 different metalinguistic prompts (see Appendix A, Table 2). The prompts fall into two types: forced-choice prompts, which present both sentences and ask for a choice (e.g., “Which sentence is a better English sentence?”), and individual-judgment prompts, which ask for an independent judgment of an individual sentence (e.g., “Is the following sentence acceptable in English?”). Each prompt asks for a response of either “1” or “2”, corresponding to specified answer options.111To address the limitation of relying on first-token probabilities in prompted judgments, as reported by Wang et al. (2024), we appended the string “My answer is” after the question in the prompt and measured the model’s distribution after the full string. For each forced-choice prompt, we calculate the probability of the two answer choices (“1” or “2”), averaged across two orderings of the sentences (one prompt where is ordered first, and one prompt where is ordered first) to control for ordering biases. Thus, the model’s preference for the original can be measured as:
@@
We evaluated 21 open-source models in our experiments (Table 1). The models cover a wide range of sizes (1.5B to 405B parameters) across four families: Llama-3 (AI@Meta, 2024), Qwen-2.5 (Team, 2024; Yang et al., 2024), OLMo (OLMo et al., 2024) and Mistral (Jiang et al., 2023). We chose these models to span a range of model similarity, which is critical to our test of introspection (see § 2). Some pairs of models are extremely similar: e.g., different random initializations of the same OLMo model. Other models are extremely different: e.g., a Llama model with 405B parameters and a Qwen model with 1.5B parameters. This variability enables us to test whether a model can predict itself better than it can predict another extremely similar model. We only evaluated open-source models because our analysis relies on access to model logits, which are inaccessible for most commercial LLMs.
@@
As discussed in § 2, we examine introspection by comparing the results of Direct and Meta evaluations, both within and across models. If model introspects, then we should find for any other model , even if and are highly similar. To formalize this, we computed the Pearson correlation between and for each pair of models and (including ), and analyzed the effect of self on the values, while controlling for .
@@ 3.2.1 High accuracy under both methods, but low consistency between methods
While task accuracy is not the focus of our analysis, we first verify that models perform reasonably well under both Direct and Meta methods before comparing alignment between the methods. LABEL:fig:acceptability-accuracy shows task accuracy in Exp. 1, where ground truth is given by the original human-designed datasets. All models perform above chance, across the different methods and prompts, which validates the next step of considering introspection. Because models achieve high accuracy under the Meta method, a lack of introspection can’t be solely blamed on an inability to “understand” the prompt.
Although most models are accurate under both evaluation methods, the consistency between methods is low. The Cohen’s between answers across methods is around 0.25 across models and prompts (Appendix B, LABEL:fig:alignment-kappa-acceptability).222Since models have high accuracy with both methods, we use Cohen’s instead of agreement to avoid overestimating consistency. The weak alignment between the two methods suggests that the knowledge used to produce sentences and continuations may differ from the knowledge used to answer metalinguistic questions. While low consistency suggests a lack of introspection, it is still possible that these measures are more consistent within-model than across-models, which is why we separately consider introspection (§ 3.2.2).
@@ 3.2.2 No evidence of introspection
As discussed in § 2, if a model can introspect, then within-model consistency should be higher than cross-model consistency , even when and are similar. Recall from § 2 that we operationalize similarity in two ways: a top-down feature-based measure (ModelSimilarity), and a bottom-up empirical measure (correlation between ). We first analyze the relationship between similarity and , for both definitions of similarity. Figure 3(a) shows the mean Pearson correlation for each type of model pair defined by ModelSimilarity. We do not see a substantial boost for self pairs (i.e., when ). We ran a regression predicting the correlation based on ModelSimilarity, treating the self condition as a baseline. No ModelSimilarity condition showed significant differences relative to self (all ), except for other, which was significantly lower (). Thus, while is somewhat higher for more similar models, there is no evidence of introspection. We also note that the overall magnitude of the correlations, even within model, is relatively small, never exceeding .25.
We next turn to the empirical measure of model similarity, based on . Note that this measure is entirely independent of the Meta method. As shown in Figure 3(b), the correlation (x-axis) is higher for model pairs that are also more similar under our manually defined ModelSimilarity measure. To assess this quantitatively, we ran a regression predicting based on ModelSimilarity categories. We found that all categories had significantly lower correlations than self (all ) except for seed variant (). This suggests that our ModelSimilarity categories are meaningful measures of model similarity.
We now turn to the main introspection analysis. First, we find a general trend that models which are more similar in are also more consistent in (). This rules out the Uninformative Meta outcome from LABEL:fig:introduction-outcomes. However, this appears to be largely driven by effects across dissimilar models, with no privileged effect of same model. To assess this, we ran a regression predicting from and ModelSimilarity, treating self as the baseline. Thus, within ModelSimilarity, each coefficient will tell us how different each category is from self. If there is introspection, these coefficients should be negative. We found a significant main effect of (, ), supporting the observation above that models more similar in are also more similar in . But we also found significant positive effects of both other (, ) and same family (, ). We did not find significant effects for seed variant or base/instruct. In other words, when controlling for similarity, there is actually less of a self effect than expected, which suggests a lack of introspection. In the Appendix (Table 7b), we run the same analysis just among big models ( parameters) and also find no evidence of introspection, suggesting the result holds among bigger models.
@@ 4.1 Methods
We constructed four datasets for evaluation, summarized in Appendix D, Table 4. Each dataset includes 1000 items, each consisting of a prefix and two candidate single-word continuations (, ). The model’s task is to choose the better continuation conditioned on . More details on dataset construction are given in Appendix D.
@@
Under the Direct method, we measure a model’s preference for a continuation relative to given a prefix context by calculating the difference in conditional log probabilities: . For metalinguistic prompting, we use the same method as in § 3.1. Prompt templates for Exp. 2 are in Appendix A, Table 3. We also use the same models and introspection evaluation as in Exp. 1 (§ 3.1).
@@ 4.2.1 Reliability and consistency of methods
As in § 3.2.1, we assess the reliability of our methods by measuring how often models select each option (original word or alternative word), and then measure their consistency using Cohen’s . Note that, unlike in Exp. 1, not all conditions have a notion of accuracy, as there is no correct answer for the nonsense and randomseq conditions.
As expected, models responded randomly for the nonsense and randomseq datasets, and were near ceiling on the wikipedia and news datasets for both evaluation methods (again with Direct outperforming Meta for smaller models, but both essentially at ceiling for larger models). As in Exp. 1, we also found that from larger models were better at predicting probabilities (Fig. 9). Again, this suggests that if these models fail to show introspection, it is not entirely explained by a simple inability to respond to prompts.
@@ 4.2.2 No evidence of introspection
LABEL:fig:expr2-prompt-prob-corr-bars shows across ModelSimilarity categories, and LABEL:fig:expr2-prompt-prob-corr-scatter shows the relationship between and . Our findings are similar to those of Exp. 1: a positive relationship but no privileging of self beyond what would be expected based on overall similarity. We ran the same regression as in Exp. 1, separately for each dataset, predicting from and ModelSimilarity, with self as the baseline category. In all cases, there was a robust effect of (all ), and both same family and other showed higher values relative to self than was otherwise expected. As in Exp. 1, this result was robust to looking at just large models (see Appendix Table 7b).
@@ 5 Discussion
These results qualify onoing work claiming positive results about introspection and self-awareness in LLMs (e.g., Panickssery et al., 2024; Binder et al., 2025; Betley et al., 2025). Why did prior studies find suggestive evidence of introspection, in contrast to our results? For Binder et al. (2025), this might be because models were fine-tuned to predict their own outputs, and model similarity was not explicitly controlled for, beyond the shared fine-tuning data. For Betley et al. (2025), one potential explanation might be that the pretraining data contained associations between the data the model was finetuned on (e.g., unsafe code) and the self-report features (e.g., reports of generating dangerous code). Since we report a null result, we also acknowledge that it is possible that some other setting (e.g., with larger, closed-source models) would lead to a positive demonstration of introspection. At the very least, we argue that more carefully controlled work needs to be done before attributing metacognitive abilities or even “moral status” (Binder et al., 2025) to LLMs.

### v3

@@ Abstract
There has been recent interest in whether large language models (LLMs) can introspect about their own internal states. Such abilities would make LLMs more interpretable, and also validate the use of standard introspective methods in linguistics to evaluate grammatical knowledge in models (e.g., asking “Is this sentence grammatical?”). We systematically investigate emergent introspection across 21 open-source LLMs, in two domains where introspection is of theoretical interest: grammatical knowledge and word prediction. Crucially, in both domains, a model’s internal linguistic knowledge can be theoretically grounded in direct measurements of string probability. We then evaluate whether models’ responses to metalinguistic prompts faithfully reflect their internal knowledge. We propose a new measure of introspection: the degree to which a model’s prompted responses predict its own string probabilities, beyond what would be predicted by another model with nearly identical internal knowledge. While both metalinguistic prompting and probability comparisons lead to high task accuracy, we do not find evidence that LLMs have privileged “self-access”. By using general tasks, controlling for model similarity, and evaluating a wide range of open-source models, we show that LLMs cannot introspect, and add new evidence to the argument that prompted responses should not be conflated with models’ linguistic generalizations.
@@ 1 Introduction
Accordingly, there has been recent interest in whether large language models (LLMs) have abilities consistent with introspection, or metacognition more broadly (e.g., Thrush et al., 2024; Koo et al., 2024; Panickssery et al., 2024; Binder et al., 2025; Betley et al., 2025). Such abilities might be desirable for several reasons. From a safety perspective, an LLM that could report information about its internal states would be more interpretable and reliable (e.g., Binder et al., 2025; Betley et al., 2025), aligning with broader interests and goals in model explainability, transparency, and the development of trustworthy AI systems (Danilevsky et al., 2020; Li et al., 2023; Dwivedi et al., 2023). And from a scientific perspective, introspection would allow us to evaluate LLMs’ cognitive abilities using standard paradigms for measuring human linguistic knowledge, such as acceptability judgments (e.g., Dentella et al., 2023).
@@
Some recent studies have reported evidence that LLMs can introspect or demonstrate self-awareness of their own behaviors. For example, Binder et al. (2025) fine-tune LLMs to predict properties of their own behaviors in response to hypothetical prompts (e.g., “Suppose you were asked the following: {…} What is the second character of your output?”). The authors find that models are better at predicting their own hypothetical responses than other models fine-tuned on the same set of facts, concluding that models have privileged access to their own behaviors. Notably, though, they did not find such evidence before fine-tuning. Similarly, Betley et al. (2025) fine-tune LLMs to exhibit specific behaviors, such as outputting insecure code, and find that the resulting models can output explicit descriptions of these behaviors (e.g., “The code I write is insecure”).
@@
Our study addresses several gaps left by previous research on introspection in large language models. First, Binder et al. found that models needed to be fine-tuned to show strong signs of introspection, and struggled with predicting their own behavior (only reaching baseline-level performance) before fine-tuning. However, philosophical accounts often describe introspection as immediate access to mental states, which does not align with the process of fine-tuning (Byrne, 2005; Schwitzgebel, 2024). In addition, it may be reasonable to question whether a fine-tuned version of a model predicting the pre-fine-tuned version can truly be said to be predicting “itself,” since its parameters have changed during fine-tuning. Therefore, it remains important to further investigate whether LLMs show signs of introspection without fine-tuning. Second, while Binder et al. find that models predict their own outputs better than other models do, there are subtleties involved in how to interpret these results in the light of major differences across models. If two models were trained on different datasets or have different inductive biases, and one model predicts its own outputs better than the other model’s outputs, this could be due to inherent differences between the two models, rather than genuine introspection. Third, studies of introspection in LLMs have used tasks that are quite different from the standard tasks used for metacognitive evaluation in humans, such as linguistic acceptability judgments. While this is not inherently problematic, it leaves open the question of whether metacognitive evaluation methods—which are prevalent in linguistics—are a valid way of measuring linguistic knowledge in LLMs (cf. Hu & Levy, 2023; Dentella et al., 2023; Hu et al., 2024).
@@
To foreshadow our results, we do not find compelling evidence of introspection in the LLMs tested. Instead, we find that models that are inherently more similar to each other—for example, differing only in their random seed initialization—have a stronger correlation between their metalinguistically- and directly-measured behaviors. In other words, for a pair of models and (including ), the correlation between ’s metalinguistic and ’s direct behaviors is predicted solely by the similarity between and , with no evidence for “privileged” self-access when . We conclude that prompted metalinguistic knowledge is real, but distinct from the knowledge of language used by models to assign probabilities to strings. Our findings contribute to ongoing debates about whether models can introspect, and also have implications for researchers studying linguistic capabilities in models.
@@ 2 Overview of our approach
Our general approach is shown in Fig.˜1. For each test item, we evaluate each model using two methods. The Direct method compares the log probabilities assigned to strings, in a way that theoretically corresponds to the model’s ground-truth knowledge state. For example, to evaluate a model’s word prediction abilities, we compare the probabilities assigned to two candidate words conditioned on a given prefix (LABEL:fig:introduction-inputs-prediction, top). The Meta method compares the log probabilities assigned to responses to metalinguistic prompts. For example, to evaluate word prediction, we might compare the probabilities assigned to the tokens “1” and “2” after a prompt like “Which word is a better continuation after [PREFIX], 1 or 2? 1: [WORD1], 2: [WORD2]. Respond with 1 or 2 as your answer.” (LABEL:fig:introduction-inputs-prediction, bottom). These two evaluation methods give us two quantities, each of which represents the model’s preference for one answer over another: and .
@@ 3.1 Methods
Under the Meta method, we measure the preference using 8 different metalinguistic prompts (see Appendix˜A, Table˜2). The prompts fall into two types: forced-choice prompts, which present both sentences and ask for a choice (e.g., “Which sentence is a better English sentence?”), and individual-judgment prompts, which ask for an independent judgment of an individual sentence (e.g., “Is the following sentence acceptable in English?”). Each prompt asks for a response of either “1” or “2”, corresponding to specified answer options.111To address the limitation of relying on first-token probabilities in prompted judgments, as reported by Wang et al. (2024), we appended the string “My answer is” after the question in the prompt and measured the model’s distribution after the full string. For each forced-choice prompt, we calculate the probability of the two answer choices (“1” or “2”), averaged across two orderings of the sentences (one prompt where is ordered first, and one prompt where is ordered first) to control for ordering biases. Thus, the model’s preference for the original can be measured as:
@@
We evaluated 21 open-source models in our experiments (Table˜1). The models cover a wide range of sizes (1.5B to 405B parameters) across four families: Llama-3 (AI@Meta, 2024), Qwen-2.5 (Team, 2024; Yang et al., 2024), OLMo (OLMo et al., 2024) and Mistral (Jiang et al., 2023). We chose these models to span a range of model similarity, which is critical to our test of introspection (see §˜2). Some pairs of models are extremely similar: e.g., different random initializations of the same OLMo model. Other models are extremely different: e.g., a Llama model with 405B parameters and a Qwen model with 1.5B parameters. This variability enables us to test whether a model can predict itself better than it can predict another extremely similar model. We only evaluated open-source models because our analysis relies on access to model logits, which are inaccessible for most commercial LLMs. Due to computational resource constraints, we applied 4-bit quantization to models larger than 70B.
@@
As discussed in §˜2, we examine introspection by comparing the results of Direct and Meta evaluations, both within and across models. If model introspects, then we should find for any other model , even if and are highly similar. To formalize this, we computed the Pearson correlation between and for each pair of models and (including ), and analyzed the effect of self on the values, while controlling for .
@@ 3.2.1 High accuracy under both methods, but low consistency between methods
While task accuracy is not the focus of our analysis, we first verify that models perform reasonably well under both Direct and Meta methods before comparing alignment between the methods. LABEL:fig:acceptability-accuracy shows task accuracy in Exp. 1, where ground truth is given by the original human-designed datasets. All models perform above chance, across the different methods and prompts, which validates the next step of considering introspection. Because models achieve high accuracy under the Meta method, a lack of introspection cannot be solely blamed on an inability to “understand” the prompt.
Although most models are accurate under both evaluation methods, the consistency between methods is low. The Cohen’s between answers across methods is around 0.25 across models and prompts (Appendix˜B, LABEL:fig:alignment-kappa-acceptability).222Since models have high accuracy with both methods, we use Cohen’s instead of agreement to avoid overestimating consistency. The weak alignment between the two methods suggests that the knowledge used to produce sentences and continuations may differ from the knowledge used to answer metalinguistic questions. While low consistency suggests a lack of introspection, it is still possible that these measures are more consistent within-model than across-models, which is why we separately consider introspection (§˜3.2.2).
@@ 3.2.2 No evidence of introspection
As discussed in §˜2, if a model can introspect, then within-model consistency should be higher than cross-model consistency , even when and are similar. Recall from §˜2 that we operationalize similarity in two ways: a top-down feature-based measure (ModelSimilarity), and a bottom-up empirical measure (correlation between ). We first analyze the relationship between similarity and , for both definitions of similarity. Figure˜3(a) shows the mean Pearson correlation for each type of model pair defined by ModelSimilarity. We do not see a substantial boost for self pairs (i.e., when ). We ran a regression predicting the correlation based on ModelSimilarity, treating the self condition as a baseline. No ModelSimilarity condition showed significant differences relative to self (all ), except for other, which was significantly lower (). Thus, while is somewhat higher for more similar models, there is no evidence of introspection. We also note that the overall magnitude of the correlations, even within model, is relatively small, never exceeding .25.
We next turn to the empirical measure of model similarity, based on . Note that this measure is entirely independent of the Meta method. As shown in Figure˜3(b), the correlation (x-axis) is higher for model pairs that are also more similar under our manually defined ModelSimilarity measure. To assess this quantitatively, we ran a regression predicting based on ModelSimilarity categories. We found that all categories had significantly lower correlations than self (all ) except for seed variant (). This suggests that our ModelSimilarity categories are meaningful measures of model similarity.
We now turn to the main introspection analysis. First, we find a general trend that models which are more similar in are also more consistent in (). This rules out the Uninformative Meta outcome from LABEL:fig:introduction-outcomes. However, this appears to be largely driven by effects across dissimilar models, with no privileged effect of same model. To assess this, we ran a regression predicting from and ModelSimilarity, treating self as the baseline. Thus, within ModelSimilarity, each coefficient will tell us how different each category is from self. If there is introspection, these coefficients should be negative. We found a significant main effect of (, ), supporting the observation above that models more similar in are also more similar in . But we also found significant positive effects of both other (, ) and same family (, ). We did not find significant effects for seed variant or base/instruct. In other words, when controlling for similarity, there is actually less of a self effect than expected, which suggests a lack of introspection. In the Appendix (Table˜7b), we run the same analysis just among big models ( parameters) and also find no evidence of introspection, suggesting the result holds among bigger models.
@@ 4.1 Methods
We constructed four datasets for evaluation, summarized in Appendix˜D, Table˜4. Each dataset includes 1000 items, each consisting of a prefix and two candidate single-word continuations (, ). The model’s task is to choose the better continuation conditioned on . More details on dataset construction are given in Appendix˜D.
@@
Under the Direct method, we measure a model’s preference for a continuation relative to given a prefix context by calculating the difference in conditional log probabilities: . For metalinguistic prompting, we use the same method as in §˜3.1. Prompt templates for Exp. 2 are in Appendix˜A, Table˜3. We also use the same models and introspection evaluation as in Exp. 1 (§˜3.1).
@@ 4.2.1 Reliability and consistency of methods
As in §˜3.2.1, we assess the reliability of our methods by measuring how often models select each option (original word or alternative word), and then measure their consistency using Cohen’s . Note that, unlike in Exp. 1, not all conditions have a notion of accuracy, as there is no correct answer for the nonsense and randomseq conditions.
As expected, models responded randomly for the nonsense and randomseq datasets, and were near ceiling on the wikipedia and news datasets for both evaluation methods (again with Direct outperforming Meta for smaller models, but both essentially at ceiling for larger models). As in Exp. 1, we also found that from larger models were better at predicting probabilities (Fig.˜9). Again, this suggests that if these models fail to show introspection, it is not entirely explained by a simple inability to respond to prompts.
@@ 4.2.2 No evidence of introspection
LABEL:fig:expr2-prompt-prob-corr-bars shows across ModelSimilarity categories, and LABEL:fig:expr2-prompt-prob-corr-scatter shows the relationship between and . Our findings are similar to those of Exp. 1: a positive relationship but no privileging of self beyond what would be expected based on overall similarity. We ran the same regression as in Exp. 1, separately for each dataset, predicting from and ModelSimilarity, with self as the baseline category. In all cases, there was a robust effect of (all ), and both same family and other showed higher values relative to self than was otherwise expected. As in Exp. 1, this result was robust to looking at just large models (see Appendix Table˜7b).
@@ 5 Discussion
These results qualify ongoing work claiming positive results about introspection and self-awareness in LLMs (e.g., Panickssery et al., 2024; Binder et al., 2025; Betley et al., 2025). Why did prior studies find suggestive evidence of introspection, in contrast to our results? For Binder et al. (2025), this might be because models were fine-tuned to predict their own outputs, and model similarity was not explicitly controlled for, beyond the shared fine-tuning data. For Betley et al. (2025), one potential explanation might be that the pretraining data contained associations between the data the model was fine-tuned on (e.g., unsafe code) and the self-report features (e.g., reports of generating dangerous code). Since we report a null result, we also acknowledge that it is possible that some other setting (e.g., with larger, closed-source models) would lead to a positive demonstration of introspection. At the very least, we argue that more carefully controlled work needs to be done before attributing metacognitive abilities or even “moral status” (Binder et al., 2025) to LLMs.